Skip to end of metadata
Go to start of metadata

Weather mainly cloudy with cold northerly breeze, rain by evening.  All systems are working well.  We tidied up around the sites, repaired a hole in the roof of the profiler site shelter and installed a vent fan to prevent overheating.  Replaced webcam since the previous one had failed due to a rain.   Put the DSM at the sodar site and the cell modem at the profiler site on UPS.  Continued surveying possible sites for the NCAS wind profiler.

This was the last day of ISS setup.  Jen and Gary have already left; Lou, John and Bill are leaving tomorrow - many thanks to all!